# The Boundary Proposal

### v1

@@ 1 Introduction
The idea of vacuum creation from nothing has a long history: Vilenkin [1, 2, 3], Hartle-Hawking [4] and Linde [5] have proposed that the creation rate of de Sitter universes can be obtained by studying appropriate gravitational instantons. In fact, the relevant instanton is simply euclidean dS space, i.e. a sphere. Alternatively, one may start from the Lorentzian path integral [6, 7]. The relevant instantons then appear in the analytic continuation as those saddle points which give the dominant contribution (see [8, 9] for a recent review and critical discussions). It was furthermore observed that, if end-of-the-world (ETW) branes exist, universes with boundaries may be created from nothing. The corresponding instanton, first introduced in [10, 11] and further studied in [12, 13, 14, 15], has the topology of a ball with a spherical ETW brane as its boundary. The geometry of the creation process arises by cutting this ball in two and gluing one of its halves to an expanding Lorentzian universe with a cylindrical boundary.
@@ 2.1 Instanton geometries and actions
In case the underlying theory of quantum gravity allows for ETW branes, further instanton geometries exist. Let us briefly argue that this is a reasonable assumption: First, ETW branes can arise from shrinking extra dimensions [17]. Second, they appear naturally in type IIA string theory in the form of O8 planes. Finally, they are predicted to exist in any quantum gravity theory by the more general Cobordism Conjecture [18]. Recent studies of ETW brane examples include [19, 20, 21, 22, 23, 24, 15, 25, 26, 27, 28].
@@
We see that $|\mathcal{S}_{bos}|$ becomes small for $T\to+\infty$ and approaches the no-boundary action $\mathcal{S}_{nb}=-8\pi^{2}M_{P}^{2}\ell_{dS}^{2}$ for $T\to$-$\infty$.
@@ 3 Creation of toric universes
As has been noticed long ago [39], a toric universe with positive cosmological constant can classically evolve from zero radius without the need of an off-shell region. A more complicated analysis is required to include quantum effects, such as fluctuations of additional fields or the Casimir energy in the calculation of the creation rate.

### v2

@@ 1 Introduction
The idea of vacuum creation from nothing has a long history: Vilenkin, Hartle-Hawking and Linde have proposed that the creation rate of de Sitter universes can be obtained by studying appropriate gravitational instantons [1, 2, 3, 4, 5] (see [6, 7] for reviews). In fact, the relevant instanton is simply euclidean dS space, i.e. a sphere. Alternatively, one may start from the Lorentzian path integral [8, 9]. The relevant instantons then appear in the analytic continuation as those saddle points which give the dominant contribution (see [10, 11] for a recent review and critical discussions). It was furthermore observed that, if end-of-the-world (ETW) branes exist, universes with boundaries may be created from nothing. The corresponding instanton, first introduced in [12, 13] and further studied in [14, 15, 16, 17], has the topology of a ball with a spherical ETW brane as its boundary. The geometry of the creation process arises by cutting this ball in two and gluing one of its halves to an expanding Lorentzian universe with a cylindrical boundary.
@@ 2.1 Instanton geometries and actions
In case the underlying theory of quantum gravity allows for ETW branes, further instanton geometries exist. Let us briefly argue that this is a reasonable assumption: First, ETW branes can arise from shrinking extra dimensions [19]. Second, they appear naturally in type IIA string theory in the form of O8 planes. Finally, they are predicted to exist in any quantum gravity theory by the more general Cobordism Conjecture [20]. Recent studies of ETW brane examples include [21, 22, 23, 24, 25, 26, 17, 27, 28, 29, 30, 31]. [^1] We note in particular that, using an explicit regularization of the 4d ETW brane tension, actions can be obtained in a universal way from the 4d EFT[17, 31].
@@
We see that $|\mathcal{S}_{bos}|$ becomes small for $T\to+\infty$ and approaches the no-boundary action $\mathcal{S}_{nb}=-8\pi^{2}M_{P}^{2}\ell_{dS}^{2}$ for $T\to$-$\infty$. The FLRW universes created in this scenario have negative spatial curvature, similarly to those in creation process of [33].
@@ 3 Creation of toric universes
As has been noticed long ago [43, 44, 45], a toric universe with positive cosmological constant can classically evolve from zero radius without the need of an off-shell region. A more complicated analysis is required to include quantum effects, such as fluctuations of additional fields or the Casimir energy in the calculation of the creation rate.